OpenStudy (anonymous):

What number completes the square? x2 + 10x + ? A. 10 B. 100 C. 20 D. 25

Directrix (directrix):

D. 25

OpenStudy (anonymous):

10x=2*x*a

Directrix (directrix):

Take 1/2 of the b coefficient 10 and then square it.

OpenStudy (anonymous):

\[(a+b)^2=a^2+2ab+b^2\]

OpenStudy (hoblos):

a^2 + 2ab+b^2 = (a+b)2 here a=x and 2ab=10 so 2xb=10x b=5 b^2=25

OpenStudy (anonymous):

(10/2)^2=25